# DiffusionNOCS: Managing Symmetry and Uncertainty in Sim2Real Multi-Modal Category-level Pose Estimation

### 20 Feb 2024 v1

@@ I INTRODUCTION
In the fields of computer vision and robotics, object pose estimation is a fundamental problem. Recently, approaches based on neural networks [1, 2, 3] have shown great progress for instance-level object pose estimation, quickly saturating the existing benchmarks [4]. Consequently, a more challenging benchmark was proposed [5], relaxing instance-level object pose estimation to category-level and evaluating the pose of unseen objects within a given category. While a significant number of methods have made progress in category-level pose estimation [6, 7, 8, 9, 10], many challenges such as handling ambiguity and generalization persist.
@@
In this work, we introduce DiffusionNOCS, a novel category-level dense pose estimation method based on diffusion models, with multi-modal input representations as shown in Fig. 1. The probabilistic nature of our method allows it to naturally handle uncertainty, improving robustness and demonstrating state-of-the-art performance on various real datasets despite being trained solely on synthetic data. The contributions of this paper are as follows:
@@
    zero-shot generalization benchmark demonstrating how existing state-of-the-art methods perform in challenging real-world environments featuring occlusions and lighting changes. Our proposed approach outperforms all synthetic baselines and despite being trained only on synthetic data, outperforms methods using real data on average.
@@ II-A Object Pose Estimation
Most existing pose estimation methods can be grouped into three categories: direct pose regression [13, 14, 15, 16] (learning to output the pose directly from data), template-matching [17, 18, 19, 20] (comparing inputs with a predefined set of templates), and correspondence-based methods [2, 1, 21] (learning to estimate correspondences between input data and corresponding 3D models of interest) - each having their own benefits and drawbacks. Overall, the current state-of-the-art instance and category-level methods [3, 1, 22, 5, 23, 10] for object pose estimation are almost exclusively correspondence-based approaches. Correspondence-based methods have clear benefits when handling occlusions and generalizing to different cameras. However, they typically require carefully curated datasets and hand-crafted operations to handle symmetrical objects to solve the ambiguity of 6D object pose. For example, Wang et al. [5] designed loss functions for specific symmetry types. This is because the deterministic nature of their training has no explicit notion of uncertainty. Since real-world objects often exhibit symmetry, this poses a challenge when scaling to larger training datasets and applying these methods in the real world.
@@ II-B Handling Symmetry and Uncertainty
While common pose estimation methods have shown great promise, they often lack the capability to provide information about uncertainty and are over-confident in their predictions [24]. In particular, they often rely on extensive prior knowledge of object geometry to address ambiguities coming from symmetry and occlusions. Such methods often require a dataset of symmetry-labeled objects [5], which is impractical in real-world applications. Another family of methods uses a probabilistic formulation that offers a means to capture uncertainty in the instance of interest directly. Kendall et al. [25] augmented PoseNet [13] with uncertainty by sampling the posterior to approximate probabilistic inference. Bishop et al. [26] learned to predict parameters of a Gaussian mixture distribution by using a neural network which in turn can be used to infer the uncertainty in a prediction based on the variance of the predicted distribution. Gilitschenski et al. [27] presented orientation uncertainty learning that use the Bingham distribution, which is a promising probabilistic approach. However, despite its effectiveness in representing axis symmetry, managing discrete symmetry is challenging as the multi-modal distribution needs to be handled and a mixture Bingham distribution requires many heuristics [28, 29]. On the other hand, diffusion models offer a promising alternative, as these models can directly handle the multi-modality that is present in the training data [30]. Recently, Zhang et al. [31] used a generative diffusion approach to estimate poses naturally handling all symmetry types. Our method also uses diffusion, but we instead estimate dense canonical maps that recover both partial geometry and pose, allowing us to not only effectively handle symmetry and occlusions, but also generalize to different camera models.
Figure 2: DiffusionNOCS Pipeline: To prepare input representations from RGB, Depth and given 2D bounding box, the operations of masking, warping, resizing are conducted sequentially. To extract low-dimensional DINOv2 features with fixed shape, PCA and resizing are applied. Then, the diffusion model is conditioned by available inputs and estimates NOCS maps from noise images. Lastly, 6D pose and confidence are estimated via the point registration between denoised NOCS maps and partial point clouds from mask, depth, and given intrinsics.
@@ III Method
Task Description. We seek to estimate the 6D pose of unknown instances in known categories from a single image. To train the model, we use only synthetic data generated via rendering openly available 3D models, resulting in respective RGB, depth images, and category IDs. At inference time, our method requires only a 2D bounding box paired with a depth or RGB image, but can flexibly accommodate more input modalities if available.
Overview. Rather than attempting to predict the pose directly from images, we instead predict dense canonical maps, a powerful intermediate representation from which object pose estimates can be extracted. We propose a dense canonical map estimator via diffusion models that does not require any special operations to handle any type of symmetry. We name our method DiffusionNOCS as the normalized object coordinate space (NOCS) [5] is leveraged as the dense canonical map, which our diffusion model is trained to predict. In this section, we first explain the network architecture and its input representations, then propose a synthetic data generation method and an entire pipeline for category-level object pose estimation.
@@ III-A NOCS Preliminaries
We utilize the NOCS as proposed in previous work [5], where each object is canonicalized such that all points exist in some space ${x,y,z}\in[0,1]$. The NOCS map is then the projection of the observed surface of the object into the camera frame, given intrinsic information. The resulting representation is an image-like quantity for which the channel dimension encodes the 3D coordinate of the surface point represented by the pixel. Given a predicted NOCS map, the pose of an object can be extracted via standard pose estimation techniques (i.e., PnP if only RGB is available and pointcloud registration if depth information is available). We observe that multiple NOCS map hypotheses can therefore result in multiple pose hypotheses.
@@ III-B NOCS Estimator with Diffusion Models
One of the key insights of this work is to leverage Denoising Diffusion Probabilistic Models (DDPMs) [32] for NOCS map estimation, enabling the prediction of multiple possible maps via multiple noises. Although DDPMs [32] are commonly used for producing photo-realistic images from noise, we make two modifications for our purpose. First, we change the output from photo-realistic images to NOCS maps. Second, we condition the denoising process on images and category IDs. As such, we use the following loss for training:
@@
where $k$ represents the denoising iteration. $\epsilon^{k}$ is a random Gaussian noise map with proper variance for iteration $k$. $N^{0}$ represents the raw NOCS maps, which are the ground truth from the dataset. $\epsilon_{\theta}$ is a noise estimation network with parameters $\theta$. $I$ and $C$ represent images and category IDs, respectively, for the conditioning of the noise estimator. For the training details, we refer the reader to DDPMs [32]. At the training phase, the discrete denoising scheduler [32] with 1000 iterations is used as a noise scheduler.
The network overview is shown in the middle part of Fig. 2. Our main network is based on 2D U-Net [33] with ResNet blocks [34]. Images to be conditioned are concatenated to the noise before being passed to the U-Net. Additionally, two more components are utilized to manage the representation of time and category IDs. The first component is a sinusoidal positional encoding [35] and an MLP for time representation. The second is an embedding layer that serves as a lookup table for the category IDs. These outputs are then concatenated and used as input for each layer of the U-Net model.
While the DDPM scheduler is utilized during training with 1000 iterations, repeating such a large number of iterations at inference time can be quite time-consuming. This is a significant concern for practical robotic applications where time efficiency is crucial. As such, during inference we utilize the DPM-solver [36], which is a fast ODE-solver for diffusion models. This allows us to significantly decrease the iterations from 1000 to 10, improving the time efficiency to enable more effective use for real-world robotics applications.
@@ III-C Representations for Conditioning
To estimate 6D poses for a large variety of objects with limited amount of real or synthetic data, the choice of input representation is critical. In our work we choose to use four types of conditioning information: surface normals, RGB, DINOv2 [37] semantic features, and category IDs.
@@
Depth information is essential for category-level pose estimation pipelines. Not only because it can help resolve scale-translation ambiguity, but also because it gives a good estimate of the partial object geometry. However, using unprocessed depth with learning based methods poses certain challenges since it entangles scale, translation and camera intrinsic parameters, requiring large amounts of data to avoid overfitting. In our work, we decouple estimation of these parameters and condition our NOCS estimator network on scale/translation invariant surface normals computed directly from input depth data.
@@
While RGB data can be highly useful, it is also a highly challenging modality. Even a single class of objects can contain vast varieties of materials resulting in a diverse set of possible object appearances, while being very similar geometrically. Moreover, material appearance can further dramatically change under different lighting conditions. As a result, it requires large amounts of data to train methods that faithfully generalize to these changes, which can be very expensive. Thankfully, powerful feature extractors trained on vast amounts of RGB data already exist and are readily available. In our work, we leverage DINOv2 [37]—a well-known foundation model trained in a self-supervised manner, which proved to be useful for a variety of image-level visual tasks.
To use DINOv2 features as a conditioning modality, we first preprocess them. The original feature dimension of the ViT-S-based DINOv2 is 384, which is much larger than all the previous modalities combined. This high-dimensional input not only makes inference and training speed slower, but also requires a larger network resulting in a larger memory footprint. To solve this problem, we follow Goodwin et al. [12] and utilize the principal component analysis (PCA) to reduce the dimensionality from N to M (N$>$M). The DINOv2 features via PCA with $N=384$, $M=3$ are visualized in Fig. 3, and show consistent coloring for different object instances from both real and synthetic images.
@@
Finally, we also include category IDs as an input modality. They are readily available from common 2D object detectors and serve as a valuable global semantic prior.
@@ III-D Selectable Input Representations
For the learning-based category-level pose estimation, depth only or RGB-depth is a common input representation choice. To make our method more flexible, we propose an approach that can freely take the available inputs at inference time with a single network without re-training. To do this, we drop each condition at a certain rate during training as follows:
@@
$I$ in Eq. 1 is separated to $I_{normal}$, $I_{RGB}$, and $I_{dino}$. When a condition is dropped, it is replaced with a null value $\emptyset$, which in practice means a tensor of the same dimension as the original condition, but filled with zeros. For example, dropping an image conditioning variable means that all pixels are filled by zeroes, and we let the zeroth ID represent an unknown category. Intuitively, even if all conditioning modalities are dropped, the model attempts to generate NOCS maps from only noise, i.e., it performs unconditional diffusion. On the other hand, if an RGB image is available, the model generates NOCS by conditioning on the RGB image as shown in Fig. 4. At inference time, available inputs can be used without re-training. We observe that this conditioning method is both powerful and flexible, enabling new useful representations to be incorporated as they are developed.
Figure 4:  Estimated NOCS maps from Different Conditions: Thanks to its generative nature, our method is capable of recovering plausible NOCS maps even when provided with only noise image input, i.e. the top left example shows a NOCS map highly resembling the camera object. Adding a category ID to the input limits the output distribution to the objects of a specified category (top right). When RGB/Normal images are provided, the silhouettes of the resulting NOCS maps are aligned with the inputs. Moreover, the output NOCS faithfully preserves geometric information provided by the surface normals as can be seen from the bottom right example where a hollow bottle is reconstructed. In contrast, the filled bottle is recovered when RGB information alone is provided (bottom left).
@@ III-E Synthetic Data Generation
To make DiffusionNOCS generalize across various real environments using just a single dataset, we introduce a dataset agnostic method of synthetic data generation for our training. We assume 3D models with aligned coordination system are given. RGB and depth images from 162 camera poses using the 3D models are rendered on white background as shown in Fig. 5. The camera poses are created via icosahedron sampling with 2 subdivision levels. Then, we create normal and ground truth NOCS maps given rendered depth as well as the known poses and sizes of objects. We estimate the DINOv2 features for all RGB images in the training dataset and apply principle component analysis (PCA) to reduce the dimension. As the resolution of DINOv2 features becomes lower than input RGB, we resize with nearest neighbor interpolation to ensure consistent dimensions.
@@
To aid in generalization, we augment the data in two ways. The first is to apply in-plane rotation for the images at the training time instead of making camera poses more various at rendering time. The second is to apply lighting augmentation based on Phong reflection model [38] with the normal maps at training time. Because of this, the lighting variation in the rendered RGB images can be less since it can be added after the rendering. At inference time, we also apply a warping operation to move the object to the center of image to ensure consistency with the training images. We show how well the generalization of DiffusionNOCS trained by the synthetic data works across various real environments in Sec. IV-B.
@@ III-F Pipeline for Category-Level Pose Estimation
The overview of entire inference pipeline for the 6D pose estimation is described in Fig. 2. For the condition of DiffusionNOCS, we aim to prepare fixed square shape normal maps, RGB images, and DINOv2 features with white background from given RGB, depth images and intrinsic parameters with 2D bounding box from detection. To create the normal maps and RGB images with fixed square shape, the cropping, warping and resizing operations are conducted sequentially on given RGB and depth images. For the cropping and warpping, detected square bounding boxes and their center are used respectively. For the creation of low-dimentional DINOv2 features, the processed RGB image are fed to DINOv2 network, then PCA and resizing to the same size of the RGB image are applied. After the NOCS map estimation via DiffusionNOCS, the square NOCS maps are reverted to the original image shape by re-warping and resizing, then a point cloud registration is conducted between points from masked depth image with intrinsic and points from NOCS maps to estimate 6D pose and scale. TEASER++, a fast point cloud registration algorithm [39], is utilized for the registration. After the registration, we treat the inlier rate from TEASER++ as the confidence of the estimated poses. This confidence is an important value for the next section.
Figure 6: Final Pose Selection: By different noise images, multiple NOCS maps can be estimated. TEASER++ predicts poses and their confidence by the NOCS maps. The final pose is selected with the highest confidence.

### 05 Mar 2024 v2

@@ I INTRODUCTION
In the fields of computer vision and robotics, object pose estimation is a fundamental problem. Recently, approaches based on neural networks [1, 2, 3] have shown great progress for instance-level object pose estimation, quickly saturating the existing benchmarks [4]. Consequently, a more challenging benchmark was proposed [5], relaxing instance-level object pose estimation to category-level and evaluating the pose of unseen objects within a given category. While a significant number of methods have made progress in category-level pose estimation [6, 7, 8], many challenges such as handling ambiguity and generalization persist.
@@
In this work, we introduce a novel category-level dense pose estimation method based on diffusion models, with multi-modal input representations as shown in Fig. 1. The multi-modal inputs and the probabilistic nature of our method allow it to naturally handle uncertainty, improving robustness and demonstrating state-of-the-art performance on various real datasets despite being trained solely on synthetic data. The contributions of this paper are as follows:
@@
    zero-shot generalization benchmark demonstrating how existing state-of-the-art methods perform in challenging real-world environments featuring pose distribution, lighting, and occlusion changes. Despite being trained only on synthetic data, our proposed approach outperforms all baselines using synthetic and real data on average.
@@ II-A Object Pose Estimation
Most existing pose estimation methods can be grouped into three categories: direct pose regression [11, 12, 13] (learning to output the pose directly from data), template-matching [14, 15, 16] (comparing inputs with a predefined set of templates), and correspondence-based methods [2, 1] (learning to estimate correspondences between input data and corresponding 3D models of interest) - each having their own benefits and drawbacks. Overall, the current state-of-the-art instance and category-level methods [3, 5, 6] for object pose estimation are almost exclusively correspondence-based approaches. Correspondence-based methods have clear benefits when handling occlusions and generalizing to different cameras. However, they typically require carefully curated datasets and hand-crafted operations to handle symmetrical objects to solve the ambiguity of 6D object pose. For example, Wang et al. [5] designed loss functions for specific symmetry types. This is because the deterministic nature of their training cannot manage multiple feasible pose hypotheses. Since real-world objects often exhibit symmetry, this poses a challenge when scaling to larger training datasets and applying these methods in the real world.
@@ II-B Handling Symmetry and Uncertainty
While common pose estimation methods have shown great promise, they often lack the capability to provide information about uncertainty and are over-confident in their predictions [17]. In particular, they often rely on extensive prior knowledge of object geometry to address ambiguities coming from symmetry and occlusions. Such methods often require a dataset of symmetry-labeled objects [5], which is impractical in real-world applications. Another family of methods uses a probabilistic formulation that offers a means to capture uncertainty in the instance of interest directly. Kendall et al. [18] augmented PoseNet [11] with uncertainty by sampling the posterior to approximate probabilistic inference. Bishop et al. [19] learned to predict parameters of a Gaussian mixture distribution by using a neural network, which in turn can be used to infer the uncertainty based on the variance of the predicted distribution. Gilitschenski et al. [20] presented orientation uncertainty learning via Bingham distribution which is suited to represent the rotational ambiguity. However, despite its effectiveness in representing axial symmetry, managing discrete symmetry is challenging as the multi-modal distribution needs to be handled [21]. On the other hand, diffusion models offer a promising alternative, as these models can directly handle the multi-modality that is present in the training data [22]. Recently, Zhang et al. [23] used a generative diffusion approach to estimate poses from partial point clouds, naturally handling all symmetry types. Our method also uses diffusion, but we instead estimate dense canonical maps from multi-modal image inputs that recover both pose and partial geometry, which not only enables us to handle symmetry and other types of ambiguity arising from object occlusions and truncation but also provides robustness against the potential degradation of noisy input data.
Figure 2: Pipeline Overview: To prepare input representations from RGB, depth and given 2D bounding box, the operations of masking, warping, resizing are conducted sequentially. For the segmentation, SAM [24] with box prompts is used. To extract low-dimensional DINOv2 features with fixed shape, PCA and resizing are applied. Then, the diffusion model is conditioned by available inputs and estimates NOCS maps from noise images. Lastly, 6D pose and confidence are estimated via the point registration between denoised NOCS maps and partial point clouds from mask, depth, and given intrinsics [25].
@@ III Method
Task Description. We seek to estimate the 6D pose of unknown instances in known categories from a single image. At inference time, we assume a depth image, 2D bounding box and category ID encompassing detected objects are given, but can flexibly accommodate more inputs, such as RGB, if available. To train the model, we use only synthetic data generated via rendering openly available 3D models, resulting in respective RGB, depth images, and category IDs.
Overview. We propose a dense canonical map estimator via diffusion models that does not require special operations to handle any type of symmetry. We name our method DiffusionNOCS as the normalized object coordinate space (NOCS) [5] is leveraged as the dense canonical map, which our diffusion model is trained to predict. In this section, we first explain the network architecture and its input representations, then propose a synthetic data generation method and an entire pipeline for category-level object pose estimation.
@@ III-A NOCS Preliminaries
We utilize the NOCS as proposed in previous work [5], where each object is canonicalized such that all points exist in some space ${x,y,z}\in[0,1]$. The NOCS map is then the projection of the observed surface of the object into the camera frame. The resulting representation is an image-like quantity for which the channel dimension encodes the 3D coordinate of the surface point represented by the pixel. Given a predicted NOCS map, the pose of an object can be extracted via standard pose estimation techniques (i.e., PnP if only RGB is available and point cloud registration if depth information is available).
@@ III-B NOCS Estimator with Diffusion Models
One of the key insights of this work is to leverage Denoising Diffusion Probabilistic Models (DDPMs) [26] for NOCS map estimation, enabling the prediction of multiple possible maps via multiple noise samples. Although DDPMs [26] are commonly used for producing photo-realistic images from noise, we make two modifications for our purpose. First, we change the output from photo-realistic images to NOCS maps. Second, we condition the denoising process on images and category IDs. As such, we use the following loss for training:
@@
where $k$ represents the denoising iteration. $\epsilon^{k}$ is a random Gaussian noise map with proper variance for iteration $k$. $N^{0}$ represents the raw NOCS maps, which are the ground truth from the dataset. $\epsilon_{\theta}$ is a noise estimation network with parameters $\theta$. $I$ and $C$ represent images and category IDs, respectively, for the conditioning of the noise estimator. For the training details, we refer reader to DDPMs [26]. At the training phase, the discrete denoising scheduler [26] with 1000 iterations is used as a noise scheduler.
The network overview is shown in the middle of Fig. 2. Our main network is based on 2D U-Net [27] with ResNet blocks [28]. Images to be conditioned are concatenated to the noise before being passed to the U-Net. Additionally, two more components are utilized to manage the representation of time and category IDs. The first component is a sinusoidal positional encoding [29] and an MLP for time representation. The second is an embedding layer that serves as a lookup table for the category IDs. These outputs are then added and used as input for each layer of the U-Net model.
While the DDPM scheduler is utilized during training with 1000 iterations, repeating such a large number of iterations at inference time can be quite time-consuming. This is a significant concern for practical robotic applications where time efficiency is crucial. As such, during inference we utilize the DPM-solver [30], which is a fast ODE-solver for diffusion models. This allows us to significantly decrease the iterations from 1000 to 10, improving the time efficiency to enable more effective use for real world robotics applications.
@@ III-C Representations for Conditioning
To estimate 6D poses for a large variety of objects with limited amount of real or synthetic data, the choice of input representation is critical. In our work, we choose to use four types of conditioning information: surface normals, RGB, DINOv2 [31] semantic features, and category IDs.
@@
Depth information is essential for category-level pose estimation pipelines. Not only because it can help resolve scale-translation ambiguities, but also because it gives a good estimate of the partial object geometry. However, using unprocessed depth with learning-based methods poses certain challenges since it entangles scale, translation and camera intrinsic parameters, requiring large amounts of data to avoid overfitting. In our work, we decouple estimation of these parameters and condition our NOCS estimator network on scale/translation invariant surface normals computed directly from input depth data.
@@
While RGB data can be highly useful, it is also a highly challenging modality. Even a single class of objects can contain a large variety of materials resulting in a diverse set of possible object appearances, while being very similar geometrically. Moreover, material appearance can further dramatically change under different lighting conditions. As a result, it requires large amounts of data to train methods that faithfully generalize to these changes, which can be very expensive. Thankfully, powerful feature extractors trained on vast amounts of RGB data already exist and are readily available. In our work, we leverage DINOv2 [31]—a foundation model trained in a self-supervised manner, which proved to be useful for a variety of image-level visual tasks [10].
To use DINOv2 features as a conditioning modality, we first preprocess them. The original feature dimension of the ViT-S-based DINOv2 is 384, which is much larger than all the previous modalities combined. This high-dimensional input not only makes inference and training speed slower, but also requires a larger network resulting in a larger memory footprint. To solve this problem, we follow Goodwin et al. [10] and utilize principal component analysis (PCA) to reduce the dimensionality from N to M (N$>$M). The DINOv2 features via PCA with $N=384$, $M=3$ are visualized in Fig. 3, and show consistent coloring for different object instances from both real and synthetic images.
@@
Category IDs are also included as an input modality. They are readily available from 2D object detectors and serve as a valuable global semantic prior.
@@ III-D Selectable Input Representations
For the learning-based category-level pose estimation, depth only or RGB-depth is a common input representation. To make our method more flexible, we propose an approach that can freely take the available inputs at inference time with a single network without re-training. To do this, we drop each condition at a certain rate during training as follows:
@@
$I$ in Eq. 1 is separated to $I_{normal}$, $I_{RGB}$, and $I_{dino}$. When a condition is dropped, it is replaced with a null value $\emptyset$, which in practice means a tensor of the same dimension as the original condition, but filled with zeros. For example, dropping an image conditioning variable means that all pixels are filled by zeroes, and we let the zeroth ID represent an unknown category. Even if all conditioning modalities are dropped, the model attempts to generate NOCS maps from only noise, i.e., it performs unconditional diffusion. On the other hand, if an RGB image is available, the model generates NOCS by conditioning on the RGB image as shown in Fig. 4. At inference time, available inputs can be used without re-training. We observe this conditioning method is both powerful and flexible, enabling new useful representations to be incorporated as they are developed.
Figure 4:  Estimated NOCS maps from Different Conditions: As a generative model, our method is capable of recovering plausible NOCS maps even when provided with only noise image input, i.e. the top left example shows a NOCS map highly resembling the ”laptop” object. Adding a category ID to the input limits the output distribution to the objects of a specified category (top right). When RGB images or surface normals are provided, the silhouettes of the resulting NOCS maps are aligned with the inputs. Moreover, the output NOCS faithfully preserves geometric information provided by the surface normals as can be seen from the bottom right example where a hollow bottle is reconstructed. In contrast, the filled bottle is recovered when RGB information alone is provided (bottom left).
@@ III-E Synthetic Data Generation
To make DiffusionNOCS generalize across various real environments using just a single dataset, we introduce a domain-agnostic synthetic data generation method. We assume 3D models with aligned coordination system are given. RGB and depth images from 162 camera poses using the 3D models are rendered on white background as shown in Fig. 5. The camera poses are created via icosahedron sampling with 2 subdivision levels. Then, we create normal and groundtruth NOCS maps given rendered depth as well as the known poses and sizes of objects. We estimate the DINOv2 features for all RGB images in the training dataset and apply principle component analysis (PCA) to reduce the dimension. As the resolution of DINOv2 features becomes lower than input RGB, we resize with nearest neighbor interpolation to ensure consistent dimensions. To aid in generalization, we augment the data in three ways. The first is to apply in-plane rotation for the images at the training time instead of making camera poses more various at rendering time. The second is to apply lighting augmentation based on Phong reflection model [32] with the normal maps at training time. Because of this, the lighting variation in the rendered RGB images can be less since it can be added after the rendering. The third is conducting Cutout [33] by random shapes that are rectangle, circle, triangle, and ellipse to improve the robustness against occlusion. At inference time, we also apply a warping operation to move the object to the center of image to ensure consistency with the training images. We show how well the generalization of DiffusionNOCS trained by the synthetic data works across various real environments in Sec. IV-B.
@@ III-F Pipeline for Category-Level Pose Estimation
The overview of entire inference pipeline for the 6D pose estimation is described in Fig. 2. For the condition of DiffusionNOCS, we aim to prepare fixed square shape normal maps, RGB images, and DINOv2 features with white background from given RGB, depth images and intrinsic parameters with 2D bounding box from detection. To create the normal maps and RGB images with fixed square shape, the cropping, warping and resizing operations are conducted sequentially on given RGB and depth images. For the cropping and warpping, detected square bounding boxes and their center are used. To generate segmentations, a pretrained SAM [24] with 2D bounding box prompts is used. For the creation of low-dimentional DINOv2 features, the processed RGB image are fed to DINOv2 network, then PCA and resizing to the same size of the RGB image are applied. After the NOCS map estimation via DiffusionNOCS, the square NOCS maps are reverted to the original image shape by re-warping and resizing, then a point cloud registration is conducted between points from masked depth image with intrinsic and points from NOCS maps to estimate 6D pose and scale. TEASER++, a fast point cloud registration algorithm [25], is utilized for the registration. After the registration, we treat the inlier rate from TEASER++ as the confidence of the estimated poses. This confidence is an important value for the next section.